Fit a poisson distribution with the following data
X 0 1 2 3 4 5
F 142 156 69 27 5 1
1
Expert's answer
2019-11-12T09:48:16-0500

Suppose we conduct a Poisson experiment, in which the average number of successes within a given region is "\lambda" . Then, the Poisson probability is:


"P(x;\lambda)={e^{-\lambda}\lambda^x \over x!}"

"\begin{array}{c:c:c:c:c:c:c:c}
 X & 0 & 1 & 2 & 3 & 4 & 5 & Total\\ \hline
 F & 142 & 156 & 69 & 27 & 5 & 1& 400
\end{array}"


"\text{mean of given distribution}={\sum FX \over \sum F}="

"={142(0)+156(1)+69(2)+27(3)+5(4)+1(5) \over 400}=1"

This is the parameter "\lambda" of the Poisson distribution ":\lambda=1"

Required Poisson distribution is


"N\cdot{e^{-\lambda}\lambda^x \over x!}, \text{where } N=\sum F=400"

"\begin{array}{c:c:c}
 x & P(x) & Theoretical\ Frequency \\ \hline
 0 & 147.15 & 147
\\
 \hline
 1 & 147.15 & 147
\\ \hline
 2 & 73.58 & 74\\
 \hline
 3 & 24.53 & 25 \\
 \hline
 4 & 6.13 & 6 \\ \hline
 5 & 1.23 & 1 \\ \hline
 & & Total=400 \\
\end{array}"